# AeroVerse: UAV-Agent Benchmark Suite for Simulating, Pre-training, Finetuning, and Evaluating Aerospace Embodied World Models

### v1

@@ Abstract
Aerospace embodied intelligence aims to empower unmanned aerial vehicles (UAVs) and other aerospace platforms to achieve autonomous perception, cognition, and action, as well as egocentric active interaction with humans and the environment. The aerospace embodied world model serves as an effective means to realize the autonomous intelligence of UAVs and represents a necessary pathway toward aerospace embodied intelligence. However, existing embodied world models primarily focus on ground-level intelligent agents in indoor scenarios, while research on UAV intelligent agents remains unexplored. To address this gap, we develop AeroSimulator, a simulation platform that encompasses $4$ realistic urban scenes for UAV flight simulation. Additionally, we construct the first large-scale real-world image-text pre-training dataset, AerialAgent-Ego10k, featuring urban drones from a first-person perspective. We also create a virtual image-text-pose alignment dataset, CyberAgent-Ego500k, to facilitate the pre-training of the aerospace embodied world model. For the first time, we clearly define $5$ downstream tasks, i.e., aerospace embodied scene awareness, spatial reasoning, navigational exploration, task planning, and motion decision, and construct corresponding instruction datasets, i.e., SkyAgent-Scene3k, SkyAgent-Reason3k, SkyAgent-Nav3k and SkyAgent-Plan3k, and SkyAgent-Act3k, for fine-tuning the aerospace embodiment world model. Simultaneously, we develop SkyAgent-Eval, the downstream task evaluation metrics based on GPT-4, to comprehensively, flexibly, and objectively assess the results, revealing the potential and limitations of 2D/\///3D visual language models in UAV-agent tasks. Furthermore, we integrate over $10$ 2D/\///3D visual-language models, $2$ pre-training datasets, $5$ finetuning datasets, more than $10$ evaluation metrics, and a simulator into the benchmark suite, i.e., AeroVerse, which will be released to the community to promote exploration and development of aerospace embodied intelligence.
Index Terms:
Aerospace Embodied Intelligence, Aerospace Embodied World Model, UAV-Agent, Visual-Language Model.
@@ I Introduction
Drones have a wide range of applications, including mountainous photovoltaic inspection, river trash detection, pedestrian traffic monitoring at intersections, electric power inspection, and forest fire rescue [1]. However, these applications often depend on manual remote control of the drones. For instance, in UAV mountain photovoltaic inspections, it is necessary to deploy professional operators who spend several hours each day inspecting multiple stations. This practice can easily lead to operator fatigue, resulting in component defects and missed inspections. Therefore, there is an urgent need for UAVs equipped with autonomous intelligence to reduce costs and enhance efficiency.
@@
Aerospace embodied intelligence refers to the specialized application of embodied intelligence within the aerospace sector, focusing on empowering unmanned platforms such as satellites, drones, and aircraft to autonomously integrate perception, cognition, and action. This integration aims to facilitate egocentric active interactions with both humans and the physical environment. Over the past year, visual-language models that encode world knowledge have rapidly advanced, driven by a wealth of high-fidelity simulators and datasets[2, 3, 4, 5, 6, 7], thereby presenting new opportunities for embodied intelligence. Numerous embodied world models[8, 9, 10, 11, 12, 13, 14, 15], have emerged, significantly enhancing the capabilities of embodied agents in perceiving their surroundings and planning tasks. Consequently, this article posits that the development of an aerospace embodied world model is a crucial strategy for achieving autonomous intelligent agents for drones and represents a necessary pathway toward advancing aerospace embodied intelligence.
@@
Figure 1: The benchmark suite for the aerospace embodiment world model, AeroVerse, comprises one simulation platform (AeroSimulator), two real-virtual pre-training datasets (AerialAgent-Ego10k and CyberAgent-Ego500k), five downstream task instruction datasets (SkyAgent-Scene3k, SkyAgent-Reason3k, SkyAgent-Nav3k and SkyAgent-Plan3k, and SkyAgent-Act3k ), and more than ten evaluation metrics (SkyAgent-Eval).
@@
However, these embodied world models primarily focus on indoor scenarios (e.g., robotic arms) or ground-based agents in outdoor environments (e.g., unmanned vehicles) [8, 9, 10, 11, 12]. There has been limited exploration of UAV embodied agents, particularly in the context of aerospace embodied world models that facilitate UAV autonomy, which is significantly constrained by the development of UAV embodied datasets. In contrast to indoor embodied intelligence datasets [9, 8], several key challenges arise in the construction of UAV embodied intelligence datasets:
@@
Lack of Definition of UAV Embodied Tasks. In recent years, research on ground-oriented agents has gained significant attention, leading to clearer definitions of downstream tasks such as indoor/outdoor navigation[3, 5], command following[2], and embodied question answering. However, UAV agents must comprehend the intrinsic correlations of four-dimensional space-time and perform actions under conditions of scene randomization and local observability of the environment. This involves aspects such as awareness, cognition, planning, and decision-making. The diversity and interdependence of these downstream tasks result in a lack of clarity in the task definitions for aerial-embodied agents.
@@
As illustrated in Figure 1, we address the gap in the UAV-agent dataset and enhance the training of aerospace embodied world models by constructing the first large-scale virtual-reality pre-training dataset alongside a high-quality instruction dataset. Specifically, the first-person, high-resolution real-world pre-training dataset of high-altitude drones, AerialAgent-Ego10k, is derived from the UrbanBIS dataset. Additionally, we develop the aligned pre-training dataset, CyberAgent-Ego500k, which includes perspective images, scene text descriptions, and drone attitudes. Furthermore, we create five downstream task instruction datasets: SkyAgent-Scene3k, SkyAgent-Reason3k, SkyAgent-Nav3k and SkyAgent-Plan3k, and SkyAgent-Act3k. These datasets are constructed using our established simulation platform, AeroSimulator, which employs Unreal Engine 4, the Microsoft AirSim drone simulator[16], and the 3D UrbanScene virtual city dataset[17]. This encompasses four real-world urban scenarios, i.e., Shanghai, Shenzhen, School, and Residence, with a coverage area ranging from $1.0\times 10^{5}$ to $3.7\times 10^{7}m^{2}$. AerialAgent-Ego10k is utilized to enhance the model’s ability to comprehend real urban scenes. CyberAgent-Ego500k is designed for virtual alignment pre-training of visual-language-posture for the Aerospace Embodied World Model, aiming to improve the model’s fundamental generalization capacity in simulated environments. This dataset contains 500K aligned UAV postures, first-person view images, and text descriptions, collected from four 3D urban environments. The collection principle prioritizes images featuring complex scenes, including buildings, roads, and trees. Furthermore, the downstream task instruction datasets are compiled by ten trained professional annotators who operated the UAV in a 3D urban environment for data acquisition and annotation. This process took a total of eight months and underwent rigorous quality checks to ensure the accuracy and reliability of the data, making it ideal for fine-tuning and evaluating the performance of downstream tasks.
@@
Furthermore, we develop a range of scientific automated evaluations, i.e., SkyAgent-Eval, for downstream tasks. Previous advancements have introduced various rubrics for text generation tasks, including BLUE [18], SPICE[19]. These methods assess text quality from relatively fixed and limited perspectives, such as semantic similarity and word matching, which can impede their customization and adaptability for evaluating downstream tasks involving UAV agents. Moreover, most existing methods depend on probabilistic statistics and do not align with human preferences. In contrast, large language models [20, 21, 22], trained using reinforcement learning with human feedback (RLHF)[23], generate responses that more accurately reflect human values and preferences. This makes them a viable alternative for evaluating text generation while mitigating the high costs associated with human evaluation[24]. Therefore, by leveraging the multifaceted capabilities of large language models, we propose an automated evaluation approach based on GPT-4 [20] for three types of downstream tasks, specifically LLM-Judge-Scene, LLM-Judge-Reason&\&&Nav, and LLM-Judge-Plan. This approach employs few-shot instruction and context learning to cater to the customized evaluation needs of various downstream tasks, thus facilitating a more comprehensive and objective assessment of their performance.
@@
(1) We construct the first large-scale real-world image-text pre-training dataset, AerialAgent-Ego10k, utilizing urban UAVs as the primary viewpoint. Additionally, we develop the virtual image-text-posture alignment dataset, CyberAgent-Ego500k, to pre-train the aerospace embodied world model, thereby enhancing the UAV agent’s ability to adapt to both real and virtual environments.
@@
(4) Extensive experiments are conducted using ten mainstream baselines to analyze their performance on the downstream instruction datasets. The experimental results reveal the potential and limitations of 2D/\///3D visual-language models in UAV-agent tasks, underscoring the necessity of constructing an aerospace embodied world model.
(5) We design over $10$ 2D/\///3D visual-language models, $2$ pre-training datasets, $5$ downstream task instruction datasets, and $10+$ evaluation metrics, as well as a simulator featuring $4$ urban scenarios, into a benchmark suite, AeroVerse, which will be released to the community to advance the field of aerospace embodied agents.
@@ II Related Work
The three-dimensional (3D) world encompasses not only horizontal and vertical dimensions but also depth, providing richer information than two-dimensional (2D) images. Depth accurately reflects fundamental aspects of the real world and enhances the ability of embodied agents to learn from and understand their 3D environment. Furthermore, textual annotations accompanying 3D visual-language datasets assist embodied agents in perceiving their surroundings and conducting spatial reasoning. However, challenges in creating 3D datasets have led to a scarcity of such resources, with only a limited number of datasets publicly available to date. For instance, the ScanQA dataset comprises $41,363$ unique Q&A pairs, accompanied by 3D object localization annotations for $800$ indoor 3D scenes [25]. The ScanRefer dataset contains $11,046$ distinct Q&A pairs for $1,613$ indoor 3D scenes [26]. The ScanNet dataset includes $1,513$ indoor scenes featuring a total of $21$ object categories [27].
@@ II-B Embodied Intelligence Datasets
The embodied world model serves as an effective approach for empowering embodied agents to interact with their environments, autonomously plan, make decisions, act, and perform tasks similar to human capabilities. Most existing embodied world models concentrate on mobile robots in indoor settings. For example, in the embodied question-and-answer task, Abhishek et al. introduce the EQA dataset, which consists of $9,000$ question-and-answer pairs across $774$ indoor rooms [28]. In the domain of embodied task planning, Mohit et al. present the ALFRED dataset, which includes $25,743$ commands and $428,322$ image-action pairs [29]. Additionally, in the realm of embodied navigation tasks, Anderson et al. propose the R2R dataset, which comprises $21,567$ navigation instructions with an average length of $29$ words [30].
@@ III Task Formulation
Input: Multi-perspective 2D images of urban 3D scenes, including $I_{t}=\{i_{t,k}\}_{k=1}^{K}$, depth map $D_{t}=\{d_{t,k}\}_{k=1}^{K}$, multi-perspective camera pose $P_{t}=\{p_{t,k}\}_{k=1}^{K}$, position $P_{history}=\{p_{n}\}_{n=0}^{N-1}$, $I_{history}=\{i_{n}\}_{n=0}^{N-1}$, $A_{history}=\{a_{n}\}_{n=0}^{N-1}$ from 00 to $N$-$1$, and target pose $p_{end}$.
@@ IV Simulation Platform
Simulator. To simulate a realistic drone flight scenario, we utilize Unreal Engine $4$ to load urban environments and select AirSim[16] for constructing the drone model. This enables us to develop a simulator, named AeroSimulator, capable of facilitating multiple action spaces for the drone, as illustrated in Figure 3. Adhering to the real-to-sim-to-real paradigm, we select four representative scenes from the high-quality UrbanScene3D dataset[17] created by Lin et al.: Shenzhen, Shanghai, School, and Residence, all derived from 3D reconstructions of actual physical locations. Furthermore, the simulator accommodates various lighting conditions (day, evening, night, etc.), seasonal variations (spring, summer, autumn, winter), and climatic modes (sunny, cloudy, light snow, etc.), thereby enhancing the transferability of the trained drone agent to real-world applications. Within the simulator, the drone can continuously navigate the urban environment we have loaded, capturing data visually through an integrated RGB, depth, and object segmentation cameras, which output corresponding first-person perspective images in real time.
Scenes. To bridge the gap between transferring drone intelligent agents from simulated environments to real-world scenarios, we utilize UrbanScene3D[17], a large-scale data platform specifically designed for urban scene perception and reconstruction. This platform comprises over 128,000 high-resolution images captured from various cities. The selected 3D scenes from four cities, as illustrated in Figure 3, feature detailed architectural elements, including office buildings, shopping centers, residential complexes, bus stations, and subway entrances and exits. Additionally, these scenes encompass specific street details such as lanes, sidewalks, crossroads, traffic signals, and road markings, along with other urban features like streetlights, signs, trees, shrubs, and lawns. These attributes facilitate the exploration of diverse urban environments by drone intelligent agents. Among the cities, Shanghai presents the most extensive urban scene, featuring $6,850$ objects and covering an area of $3,700$ hectares. This extensive environment is advantageous for training UAV agents in long-distance navigation and path planning. In contrast, the urban scene in Shenzhen is relatively compact, covering an area of $300$ hectares with only $1,126$ objects; however, it enhances the spatial reasoning capabilities of drone intelligent agents in smaller settings. Furthermore, the campus area, which spans $130$ hectares and contains $178$ objects, and the residential zone, covering $30$ hectares with $34$ objects, focus on localized environments characterized by dense buildings and obstacles such as trees and equipment. This concentration improves scene understanding and decision-making skills, including obstacle avoidance.
@@
RGB Camera. Captures RGB images with a resolution of $1920$$\times$$1080$, saved in PNG format.

### v2

@@ Abstract
Aerospace embodied intelligence aims to empower unmanned aerial vehicles (UAVs) and other aerospace platforms to achieve autonomous perception, cognition, and action, as well as egocentric active interaction with humans and the environment. The aerospace embodied world model serves as an effective means to realize the autonomous intelligence of UAVs and represents a necessary pathway toward aerospace embodied intelligence. [Background] However, existing embodied world models primarily focus on ground-level intelligent agents in indoor scenarios, while research on UAV intelligent agents remains unexplored, lacking systematic and standardized benchmark suites. [Aim] To address this gap, this study aims to construct a comprehensive benchmark suite, AeroVerse, to facilitate the simulation, pre-training, finetuning, and evaluation of aerospace embodied world models. [Innovations] We develop AeroSimulator, a simulation platform that encompasses four realistic urban scenes for UAV flight simulation. Additionally, we construct the first large-scale real-world image-text pre-training dataset from a first-person UAV perspective, AerialAgent-Ego15k, and create a virtual image-text-pose alignment dataset, CyberAgent-Ego500k, to facilitate the pre-training of the aerospace embodied world model. We clearly define five downstream tasks for the first time, i.e., aerospace embodied scene awareness, spatial reasoning, navigational exploration, task planning, and motion decision, and have constructed corresponding instruction datasets for fine-tuning. We also develop SkyAgent-Eval, a downstream task evaluation system based on GPT-4. Furthermore, we propose SkyAgentX, the first UAV-agent large model integrating “perception-reasoning-navigating-planning”, which innovatively incorporates aerospace embodied chain-of-thought mechanism and multitask curriculum learning strategy. [Results] By benchmarking ten mainstream models, our results reveal the significant limitations of existing 2D/3D visual-language models in complex aerospace embodied tasks and demonstrate the superior performance of SkyAgentX, which outperforms existing methods by an average of 8.52% across four core tasks, underscoring the necessity and contribution of our work. The AeroVerse benchmark suite will be released to the community to promote exploration and development of aerospace embodied intelligence. (https://github.com/06f081zyd/AeroVerse)
@@ I Introduction
Drones have a wide range of applications, including mountainous photovoltaic inspection, river trash detection, pedestrian traffic monitoring at intersections, electric power inspection, and forest fire rescue [Dji1]. However, these applications often depend on manual remote control of the drones. For instance, in UAV mountain photovoltaic inspections, it is necessary to deploy professional operators who spend several hours each day inspecting multiple stations. This practice can easily lead to operator fatigue, resulting in component defects and missed inspections. Therefore, there is an urgent need for UAVs equipped with autonomous intelligence to reduce costs and enhance efficiency.
@@
Aerospace embodied intelligence refers to the specialized application of embodied intelligence within the aerospace sector, focusing on empowering unmanned platforms such as satellites, drones, and aircraft to autonomously integrate perception, cognition, and action. This integration aims to facilitate egocentric active interactions with both humans and the physical environment. Over the past year, visual-language models that encode world knowledge have rapidly advanced, driven by a wealth of high-fidelity simulators and datasets[Shridhar2019ALFREDAB, ku-etal-2020-room, Li2021iGibson2O, Qi2019REVERIERE, Shen2020iGibson1A, Xia2019InteractiveGB, gao2025openfly], thereby presenting new opportunities for embodied intelligence. Numerous embodied world models[huang2023embodied, Hong20233DLLMIT, Driess2023PaLMEAE, Brohan2023RT2VM, Mu2023EmbodiedGPTVP, X2-VLM, EnhancingVisualGrounding, Vision-LanguageModelsASurvey], have emerged, significantly enhancing the capabilities of embodied agents in perceiving their surroundings and planning tasks. Consequently, this article posits that the development of an aerospace embodied world model is a crucial strategy for achieving autonomous intelligent agents for drones and represents a necessary pathway toward advancing aerospace embodied intelligence.
@@
Figure 1: The benchmark suite for the aerospace embodiment world model, AeroVerse, comprises one simulation platform (AeroSimulator), two real-virtual pre-training datasets (AerialAgent-Ego15k and CyberAgent-Ego500k), five downstream task instruction datasets (SkyAgent-Scene3k, SkyAgent-Reason3k, SkyAgent-Nav3k and SkyAgent-Plan3k, and SkyAgent-Act3k ), and more than ten evaluation metrics (SkyAgent-Eval).
@@
However, these embodied world models primarily focus on indoor scenarios (e.g., robotic arms) or ground-based agents in outdoor environments (e.g., unmanned vehicles) [huang2023embodied, Hong20233DLLMIT, Driess2023PaLMEAE, Brohan2023RT2VM, Mu2023EmbodiedGPTVP]. There has been limited exploration of UAV embodied agents, particularly in the context of aerospace embodied world models that facilitate UAV autonomy, which is significantly constrained by the development of UAV embodied datasets. In contrast to indoor embodied intelligence datasets [Hong20233DLLMIT, huang2023embodied], several key challenges arise in the construction of UAV embodied intelligence datasets:
@@
Lack of Definition of UAV Embodied Tasks. In recent years, research on ground-oriented agents has gained significant attention, leading to clearer definitions of downstream tasks such as indoor/outdoor navigation[ku-etal-2020-room, Qi2019REVERIERE], command following[Shridhar2019ALFREDAB], and embodied question answering. However, UAV agents must comprehend the intrinsic correlations of four-dimensional space-time and perform actions under conditions of scene randomization and local observability of the environment. This involves aspects such as awareness, cognition, planning, and decision-making. The diversity and interdependence of these downstream tasks result in a lack of clarity in the task definitions for aerial-embodied agents.
@@
As illustrated in Figure 1, we address the gap in the UAV-agent dataset and enhance the training of aerospace embodied world models by constructing the first large-scale virtual-reality pre-training dataset alongside a high-quality instruction dataset. Specifically, the first-person, high-resolution real-world pre-training dataset of high-altitude drones, AerialAgent-Ego15k, is derived from the UrbanBIS dataset. Additionally, we develop the aligned pre-training dataset, CyberAgent-Ego500k, which includes perspective images, scene text descriptions, and drone attitudes. Furthermore, we create five downstream task instruction datasets: SkyAgent-Scene3k, SkyAgent-Reason3k, SkyAgent-Nav3k and SkyAgent-Plan3k, and SkyAgent-Act3k. These datasets are constructed using our established simulation platform, AeroSimulator, which employs Unreal Engine 4, the Microsoft AirSim drone simulator[airsim2017fsr], and the 3D UrbanScene virtual city dataset[Liu2021UrbanScene3DAL]. This encompasses four real-world urban scenarios, i.e., Shanghai, Shenzhen, School, and Residence, with a coverage area ranging from $1.0\times 10^{5}$ to $3.7\times 10^{7}m^{2}$. AerialAgent-Ego15k is utilized to enhance the model’s ability to comprehend real urban scenes. CyberAgent-Ego500k is designed for virtual alignment pre-training of visual-language-posture for the Aerospace Embodied World Model, aiming to improve the model’s fundamental generalization capacity in simulated environments. This dataset contains 500K aligned UAV postures, first-person view images, and text descriptions, collected from four 3D urban environments. The collection principle prioritizes images featuring complex scenes, including buildings, roads, and trees. Furthermore, the downstream task instruction datasets are compiled by ten trained professional annotators who operated the UAV in a 3D urban environment for data acquisition and annotation. This process took a total of eight months and underwent rigorous quality checks to ensure the accuracy and reliability of the data, making it ideal for fine-tuning and evaluating the performance of downstream tasks.
@@
Furthermore, we develop a range of scientific automated evaluations, i.e., SkyAgent-Eval, for downstream tasks. Previous advancements have introduced various rubrics for text generation tasks, including BLUE [Papineni2002BleuAM], SPICE[Anderson2016SPICESP]. These methods assess text quality from relatively fixed and limited perspectives, such as semantic similarity and word matching, which can impede their customization and adaptability for evaluating downstream tasks involving UAV agents. Moreover, most existing methods depend on probabilistic statistics and do not align with human preferences. In contrast, large language models [Achiam2023GPT4TR, Touvron2023LLaMAOA, Bai2023QwenTR], trained using reinforcement learning with human feedback (RLHF)[lambert2022illustrating], generate responses that more accurately reflect human values and preferences. This makes them a viable alternative for evaluating text generation while mitigating the high costs associated with human evaluation[Zheng2023JudgingLW]. Therefore, by leveraging the multifaceted capabilities of large language models, we propose an automated evaluation approach based on GPT-4 [Achiam2023GPT4TR] for three types of downstream tasks, specifically LLM-Judge-Scene, LLM-Judge-Reason&\&Nav, and LLM-Judge-Plan. This approach employs few-shot instruction and context learning to cater to the customized evaluation needs of various downstream tasks, thus facilitating a more comprehensive and objective assessment of their performance.
Based on the high-quality datasets and evaluation metrics, we propose SkyAgentX, the first UAV-agent embodied large model integrating “perception-reasoning-navigating-planning”. Through the innovative introduction of the aerospace embodied chain-of-thought mechanism, SkyAgentX achieves closed-loop collaboration of environmental perception, spatial relationship reasoning, and multi-step task planning. Meanwhile, combined with a multitask curriculum learning strategy, the model can progressively adapt to the demands of dynamic scenarios from simple to complex. Experiments demonstrate that SkyAgentX outperforms existing methods by an average of 8.52% in four core tasks (scene perception, spatial reasoning, navigation exploration, and task planning), providing a generalized solution for autonomous drone intelligence systems and opening a new research direction for aerospace embodied intelligence.
@@
(1) We construct the first large-scale real-world image-text pre-training dataset, AerialAgent-Ego15k, utilizing urban UAVs as the primary viewpoint. Additionally, we develop the virtual image-text-posture alignment dataset, CyberAgent-Ego500k, to pre-train the aerospace embodied world model, thereby enhancing the UAV agent’s ability to adapt to both real and virtual environments.
@@
(4) SkyAgentX pioneers as the first UAV-agent embodied large model with its end-to-end “perception-reasoning-navigating-planning” framework. This is achieved by integrating aerospace embodied chain-of-thought and multitask curriculum learning, resulting in comprehensive superiority over existing methods—marked by an average performance improvement of 8.52% across four core tasks.
(5) Extensive experiments are conducted using ten mainstream baselines to analyze their performance on the downstream instruction datasets. The experimental results reveal the potential and limitations of 2D/\//3D visual-language models in UAV-agent tasks, underscoring the necessity of constructing an aerospace embodied world model.
(6) We design over $10$ 2D/\//3D visual-language models, $2$ pre-training datasets, $5$ downstream task instruction datasets, and $10+$ evaluation metrics, as well as a simulator featuring $4$ urban scenarios, into a benchmark suite, AeroVerse, which will be released to the community to advance the field of aerospace embodied agents.
@@ II Related Work
The three-dimensional (3D) world encompasses not only horizontal and vertical dimensions but also depth, providing richer information than two-dimensional (2D) images. Depth accurately reflects fundamental aspects of the real world and enhances the ability of embodied agents to learn from and understand their 3D environment. Furthermore, textual annotations accompanying 3D visual-language datasets assist embodied agents in perceiving their surroundings and conducting spatial reasoning. However, challenges in creating 3D datasets have led to a scarcity of such resources, with only a limited number of datasets publicly available to date. For instance, the ScanQA dataset comprises $41,363$ unique Q&A pairs, accompanied by 3D object localization annotations for $800$ indoor 3D scenes [2021ScanQA]. The ScanRefer dataset contains $11,046$ distinct Q&A pairs for $1,613$ indoor 3D scenes [2020ScanRefer]. The ScanNet dataset includes $1,513$ indoor scenes featuring a total of $21$ object categories [J2021ScanNet].
@@ II-B Embodied Intelligence Datasets
The embodied world model serves as an effective approach for empowering embodied agents to interact with their environments, autonomously plan, make decisions, act, and perform tasks similar to human capabilities. Most existing embodied world models concentrate on mobile robots in indoor settings. For example, in the embodied question-and-answer task, Abhishek et al. introduce the EQA dataset, which consists of $9,000$ question-and-answer pairs across $774$ indoor rooms [2018Embodied]. In the domain of embodied task planning, Mohit et al. present the ALFRED dataset, which includes $25,743$ commands and $428,322$ image-action pairs [2019ALFRED]. Additionally, in the realm of embodied navigation tasks, Anderson et al. propose the R2R dataset, which comprises $21,567$ navigation instructions with an average length of $29$ words [2018Vision].
@@ III Task Formulation
Input: Multi-perspective 2D images of urban 3D scenes, including $I_{t}=\{i_{t,k}\}_{k=1}^{K}$, depth map $D_{t}=\{d_{t,k}\}_{k=1}^{K}$, multi-perspective camera pose $P_{t}=\{p_{t,k}\}_{k=1}^{K}$, position $P_{history}=\{p_{n}\}_{n=0}^{N-1}$, $I_{history}=\{i_{n}\}_{n=0}^{N-1}$, $A_{history}=\{a_{n}\}_{n=0}^{N-1}$ from 0 to NN-11, and target pose $p_{end}$.
@@ IV Simulation Platform
Simulator. To simulate a realistic drone flight scenario, we utilize Unreal Engine $4$ to load urban environments and select AirSim[airsim2017fsr] for constructing the drone model. This enables us to develop a simulator, named AeroSimulator, capable of facilitating multiple action spaces for the drone, as illustrated in Figure 3. Adhering to the real-to-sim-to-real paradigm, we select four representative scenes from the high-quality UrbanScene3D dataset[Liu2021UrbanScene3DAL] created by Lin et al.: Shenzhen, Shanghai, School, and Residence, all derived from 3D reconstructions of actual physical locations. Furthermore, the simulator accommodates various lighting conditions (day, evening, night, etc.), seasonal variations (spring, summer, autumn, winter), and climatic modes (sunny, cloudy, light snow, etc.), thereby enhancing the transferability of the trained drone agent to real-world applications. Within the simulator, the drone can continuously navigate the urban environment we have loaded, capturing data visually through an integrated RGB, depth, and object segmentation cameras, which output corresponding first-person perspective images in real time.
Scenes. To bridge the gap between transferring drone intelligent agents from simulated environments to real-world scenarios, we utilize UrbanScene3D[Liu2021UrbanScene3DAL], a large-scale data platform specifically designed for urban scene perception and reconstruction. This platform comprises over 128,000 high-resolution images captured from various cities. The selected 3D scenes from four cities, as illustrated in Figure 3, feature detailed architectural elements, including office buildings, shopping centers, residential complexes, bus stations, and subway entrances and exits. Additionally, these scenes encompass specific street details such as lanes, sidewalks, crossroads, traffic signals, and road markings, along with other urban features like streetlights, signs, trees, shrubs, and lawns. These attributes facilitate the exploration of diverse urban environments by drone intelligent agents. Among the cities, Shanghai presents the most extensive urban scene, featuring $6,850$ objects and covering an area of $3,700$ hectares. This extensive environment is advantageous for training UAV agents in long-distance navigation and path planning. In contrast, the urban scene in Shenzhen is relatively compact, covering an area of $300$ hectares with only $1,126$ objects; however, it enhances the spatial reasoning capabilities of drone intelligent agents in smaller settings. Furthermore, the campus area, which spans $130$ hectares and contains $178$ objects, and the residential zone, covering $30$ hectares with $34$ objects, focus on localized environments characterized by dense buildings and obstacles such as trees and equipment. This concentration improves scene understanding and decision-making skills, including obstacle avoidance.
@@
[…]